the effective mass of electron in copper is m∗=1.01 me(me = mass of an electron) and the electric conductivity of copper is 5.76 ×107 Ω−1 m−1 with EF=7 eV then calculate the Fermi mean free path for copper.
1
Expert's answer
2021-09-23T08:31:31-0400

The number density of electrons:


"n=\frac{N_A}{\Mu}\rho=8.43·10^{28}\text{ m}^{-3}."

Find Fermi velocity:


"v_F=\sqrt{\frac{2E_F}{m^*}}."

Find relaxation time:


"\tau=\frac{m^*\sigma}{e^2n}."

Thus, the Fermi mean free path "\lambda" and conductivity are connected within the following equation:


"\lambda=\tau v_F=\frac{m^*\sigma}{e^2n}\sqrt{\frac{2E_F}{m^*}}=\\\space\\
=\frac{\sigma\sqrt{2m^*E_F}}{ne^2}=\\\space\\=\frac{(5.76·10^7)\sqrt{2(1.01·9.109·10^{-31})(7·1.602·10^{-19})}}{(8.43·10^{28})(1.602·10^{-19})^2}=\\\space\\
=3.82·10^{-8}\text{ ms}^{-1}."